OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). y=4/3x+8 y=4/3x+3 y=-3/4x+8 y=-3/4x+3

OpenStudy (anonymous):

what does parallel tell you about the slope of the line you are trying to figure out?

OpenStudy (anonymous):

That they have the same slope.

OpenStudy (anonymous):

exactly, so now you know the slope which is?

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

rearrange the equation of the given line into the standard form, y=mx+b

OpenStudy (anonymous):

where m is the slope

OpenStudy (anonymous):

?

OpenStudy (anonymous):

4x − 3y = −12 becomes -3y=-12-4x now divide by -3 what do you get?

OpenStudy (anonymous):

Divide what by -3?

OpenStudy (anonymous):

divide both sides by -3... so divide the -3y by -3, and the -12-4x by -3

OpenStudy (anonymous):

what will you be left with?

OpenStudy (anonymous):

1 and I don't know the other one.